If I drive 4 hours in the morning and cover 230 km, I then stop for lunch, and after lunch I drive another 3 hours and cover 155 km, what would my average speed be for the journey that day?
Answer by TimothyLamb(4379)

the true average speed for the entire journey would have to include the zero speed during the lunch stop, but the problem doesn't say how much time was spent at the lunch stop.
---
let's assume the lunch stop was 30 minutes:
---
total trip time:
t = 4 + 0.5 + 3 = 7.5 hours
---
total trip distance:
d = 230 + 155 = 385 km
---
s = d/t
s = 385/7.5
s = 51.333333333 kph
---
now, if we only want the average speed for the driving portions of the journey:
---
total trip time:
t = 4 + 3 = 7 hours
---
s = d/t
s = 385/7
s = 55 kph
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---